Jo and Blake have a combined age of 62.
Blake is 12 years older than Jo.
How old is Jo?

Found 3 solutions by Cromlix, ikleyn, MathTherapy:
Answer by Cromlix(4381)

Hi there,
Jo's age + Blake's age = 62
Blake is 12 years older than Jo
62 - 12 = 50
Jo is 50 years old
Blake is 62 years old.
Hope this helps :-)

Answer by ikleyn(52802)

.

Jo + Blake = 62


J + (J+12) = 62


2J + 12 = 62


2J      = 62 - 12 = 50


J                 = 50/2 = 25.


ANSWER.  Jo is 25 years old.

Solved (correctly).


Answer by MathTherapy(10552)


Jo and Blake have a combined age of 62.
Blake is 12 years older than Jo.
How old is Jo?
No, @CROMLIX. That won't help because you're wrong! For you to help, you need to provide correct answers to these people who ask for help!
You should know that!!